by the following equation The annual demand for cigarettes in a certain state is given QD 100,000 - 2,000 P where P is the price per gallon and Qp is packs of cigarettes demanded per year. The supply of cigarettes is given by the equation Qs 3,000 P Suppose that a 1 TL per pack tax is levied on the price of cigarettes received by sellers Calculate the efficiency loss. What portion of the tax is shifted on the consumers?
Answer #1

&s8000P 4 QD = 100, 000-2000 p Qs = 3000 (P-1) う 100,000-2000P= Sooof-3000 6300 (20-6-1 = 58,800 Pb 20.6 (en)2。 Ps 19.6 6, 62